OpenStudy (anonymous):

I need some help with my homework If 8p = -2X2 - 100, which of the following could X possibly be?

OpenStudy (anonymous):

It's x squared btw and I will award medals

OpenStudy (anonymous):

is it supposed to be 8p or80

OpenStudy (anonymous):

8p

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

idk then sorry

OpenStudy (anonymous):

It's alright thanks for trying !!!

OpenStudy (anonymous):

then medal???

OpenStudy (anonymous):

The answer options aren't one number for x

OpenStudy (anonymous):

they are like formulas

OpenStudy (anonymous):

jk i didn't do anything

OpenStudy (anonymous):

yes of course!!! you tried your best

OpenStudy (anonymous):

how do I give a medal exactly btw LOL

OpenStudy (anonymous):

that will count as answering a question

OpenStudy (anonymous):

hit best response

hero (hero):

Where is the list of choices?

OpenStudy (anonymous):

I'll post them, hold on

OpenStudy (anonymous):

you don't need to give the medal

OpenStudy (anonymous):

X = square root of 8 times p plus 100 over 2 X = square root of negative 8 times p plus 100 over negative 2 X = square root of negative 8 times p minus 100 over 2 X = square root of 8 times p minus 100 over negative 2

hero (hero):

Isolate X

OpenStudy (anonymous):

well if @Hero can walk me through it, then I probably will but I really do appreciate the fact that you tried to help out

OpenStudy (anonymous):

ok so I find the square root of x squared and get the square root of x correct?

OpenStudy (anonymous):

that is the first step in getting x by itself I think

hero (hero):

You have: 8p = -2X2 - 100 So first add 100 to both sides.

OpenStudy (anonymous):

ok so you eliminate the 100 because it is the inverse property by adding a positive to a negative so then I will get 8p+100=-2x^2 correct or naw?

hero (hero):

Correct, now divide both sides by -2

OpenStudy (anonymous):

ok so I'll get -4p+50=x^2?

OpenStudy (anonymous):

I meant -4p-50

hero (hero):

Now square root both sides.

OpenStudy (anonymous):

so I would get the square root of x = the square root of -4p-50? what next?

hero (hero):

Are you sure the original question is correct as written?

OpenStudy (anonymous):

I am ! it is 8 times p = negative 2 times x squared minus one hundred

OpenStudy (anonymous):

sometimes it is better just to write it as a verbal expression so there are no misunderstandings

hero (hero):

There's no misunderstanding except for the incorrect answer choices.

OpenStudy (anonymous):

haha no joke!!! thank you so much for your help and time I really really appreciate it !

hero (hero):

yw